Regular union employees earn $120/day, and union stewards earn $144/day. a company has a need for a total of 50 employees and has a daily payroll budget of $6072. how many of each employee will meet the company's needs and consume all the budget?
Answer by mananth(16949)

Regular employee x
Union Steward ----------- y
x+y= 50 .............1
120x+144y=6072 .............2
Eliminate y
multiply (1)by -144
-144x-144y=-7200
120x+144y=6072
Add the two equations
-24x= -1128
/24
x=47 Regular employee
plug value of x in (1)
x+y=50
47+y=50
y=50-47
y=3 Union Steward